# DART: A Diffusion-Based Autoregressive Motion Model for Real-Time Text-Driven Motion Control

### v1

@@ D.2 Scheduled Training
11:         $z_{0}^{i}=\mathcal{E}(\mathbf{H},\mathbf{X}^{i})$triangleright compress motion primitive into latent space
@@
13:         $z_{t}^{i}\leftarrow\textsc{forward\_diffusion}(z_{0}^{i},t)$
14:         $\hat{z}_{0}^{i}=\mathcal{G}_{\Theta}(z_{t}^{i},t,\mathbf{H},c^{i}$) $\triangleright$ latent denoising model prediction
15:         $\hat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\hat{z}_{0}^{i})$triangleright decode predicted latent variable to future motion frames
16:         $\nabla\leftarrow\nabla_{\Theta}\mathcal{L}(z_{0}^{i},\hat{z}_{0}^{i},\mathbf{H},\mathbf{X}^{i},\hat{\mathbf{X}}^{i})$triangleright model parameter gradient calculation
@@
22:              $z_{T}^{i}\leftarrow\textsc{forward\_diffusion}(z_{0}^{i},T)$triangleright maximum noising simulating inference time
23:              $\hat{z}_{0}^{i}=\textsc{ddpm\_sample}(\mathcal{G}_{\Theta},z_{T}^{i},T,\mathbf{H},c^{i})$triangleright full DDPM denoising loop
24:              $\hat{\mathbf{X}}=sg(\mathcal{D}(\mathbf{H},\hat{z}_{0}^{i}))$triangleright decode predicted latent variable and stop gradient
@@
We use $w_{rec}=1$ and $w_{aux}=10000$ in our experiments. The denoiser model is trained using an AdamW optimizer. The learning rate is set to $1e^{-4}$ with linear annealing. Our denoiser model is trained with scheduled training (Ling et al., 2020; Bengio et al., 2015), consisting of a fully supervised stage of 100K iterations, a scheduled stage of 100K iterations, and a rollout stage of 100K iterations. We set the maximum number of rollouts as 4. With the scheduled training, our latent motion primitive model can stably generate long motion sequences and better respond to the text prompt control even at poses that are not paired with the text prompt in the dataset.
@@ E.1 Experiment details
We apply DART to conduct online motion generation using the rollout algorithm 1 with a default seed motion $\mathbf{H}_{seed}$ of rest standing and using a classifier-free guidance of $w=5$. We use the released checkpoints of FlowMDM (Barquero et al., 2024), TEACH (Athanasiou et al., 2022), and DoubleTake (Shafir et al., 2024) for baseline comparison. We adjust the handshake size and blending length of DoubleTake to be compatible with the shortest segment length of 15 frames. The history-conditioned modification of T2M-GPT is retrained on the BABEL dataset using the original hyperparameters. At generation time, the last frames of the previous action segment are encoded into tokens and provided as the first tokens when generating the next action segment to provide history conditioning of the previous action.
@@
E.2 Perceptual User Studies.
We conduct user studies to quantitatively compare our method DART with baselines to provide a more comprehensive and convincing evaluation. We run binary user studies on Amazon Mechanical Turk(AMT) to evaluate generation realism and text-motion semantic alignment. Participants are given generation results from two different methods and are asked to select the generation that is perceptually more realistic or better aligns with the action text stream in subtitles, as illustrated in Fig. 4. During the realism evaluation, action descriptions were not displayed to eliminate distractions. We sample 256 sequences of action texts and durations from the BABEL dataset and use each method to generate motions given the action timelines. Participants are shown random pairs of results from our method and a baseline, and they are asked to choose the better generation. Each comparison is voted by 3 independent participants, with the video pairs randomly shuffled to ensure that participants do not know the method source.
Figure 4: Illustration of the binary user study interface for evaluating motion-text semantic alignment (top) and perceptual realism(bottom). Participants are requested to select the generation that is perceptually more realistic or better aligns with the action descriptions in subtitles (only visible in semantic user study).
@@ Appendix F Latent diffusion noise optimization-based control
Optimization details. We use the Adam (Kingma & Ba, 2017) optimizer with a learning rate of 0.05 to optimize latent noises $Z_{T}$ for 100 to 300 steps. The learning rate is linearly annealed to 0 and the gradient is normalized to stabilize optimization. The optimization running time is dependent on both the motion sequence length and the number of optimization steps. An example experiment of a target motion sequence of 60 frames and 100 optimization steps costs around 74 seconds.
Human-scene interaction. In human-scene interaction synthesis, we use two scene constraints to avoid human-scene collision and foot-floating artifacts. To reduce human-scene interpenetration, $cons\_coll(\cdot)$ penalizes joints colliding into scenes with negative SDF value and can be formulated as:
$cons\_coll(\mathbf{M})=-\sum_{J}(\Psi(J(\mathbf{M}))-J(\mathbf{M})_{thresh})_{-}$ (11)
where $\Psi(\cdot)$ is the signed distance function returning the signed distance from the query point to the closest scene point. $J(\cdot)$ extracts the joint locations from the motion $\mathbf{M}$. $(\cdot)_{thresh}$ is the joint-dependent contact threshold value, which is determined by the joint-skin distance in the rest pose. $(\cdot)_{-}$ denotes clipping positive values.
To reduce the occurrence of foot-floating artifacts, $cons\_cont(\cdot)$ encourages the foot to be in contact with the scene and is defined as:
$cons\_cont(\mathbf{M})=\sum_{F}(\Psi(F(\mathbf{M}))-F(\mathbf{M})_{thresh})_{+}$ (12)
where $F(\cdot)$ extracts the lower foot joint location. $(\cdot)_{+}$ denotes clipping negative values.
@@ Appendix G Reinforcement Learning-based Control
Architecture. Both the actor and critic networks are 4-layer MLPs with residual connections and a hidden dimension of 512. We apply the tanh scaling: $x=4\cdot tanh(x)$ (Ling et al., 2020) to the actor output to clip the action prediction in the range of $[-4,4]$, avoiding unbounded action predictions. The actor networks are initialized with close to zero weights to boost policy training following (Andrychowicz et al., 2020). The observation input contains a 512D text embedding, a 552D of history motion, a 1D observation of floor height relative to the human pelvis, a 1D observation of the floor plane distance from the pelvis to the goal location, and a 3D unit vector of the goal direction in the human-centric coordinates frames. We clamp the goal distance observation with a maximum value of 5m and the goal direction in the range of a 120-degree field of view to simulate egocentric human perception.
Rewards. The text-conditioned goal-reaching control policies are trained to maximize the discounted expectation of the following rewards (Zhang & Tang, 2022; Li et al., 2024; Zhao et al., 2023):
@@
$r_{ori}=\frac{\langle p^{i}-p^{i-1},g-p^{i-1}\rangle+1}{2},$ (15)
The moving orientation reward encourages the moving orientation to be aligned with the goal orientation, where $p^{i}$ is the human pelvis location at step $i$, and $g$ is the goal location.
@@
The rewards are weighted with $w_{dist}=1$, $w_{succ}=1$, $w_{ori}=0.1$, $w_{floor}=100$, $w_{skate}=10$ for the hopping controller and 100 for the others.
Appendix H Computing resources

### v2

@@ D.2 Scheduled Training
11:         $\mathbf{z}_{0}^{i}=\mathcal{E}(\mathbf{H},\mathbf{X}^{i})$triangleright compress motion primitive into latent space
@@
13:         $\mathbf{z}_{t}^{i}\leftarrow\textsc{forward\_diffusion}(\mathbf{z}_{0}^{i},t)$
14:         $\hat{\mathbf{z}}_{0}^{i}=\mathcal{G}_{\theta}(\mathbf{z}_{t}^{i},t,\mathbf{H},c^{i}$) $\triangleright$ latent denoising model prediction
15:         $\hat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\hat{\mathbf{z}}_{0}^{i})$triangleright decode predicted latent variable to future motion frames
16:         $\nabla\leftarrow\nabla_{\theta}\mathcal{L}(\mathbf{z}_{0}^{i},\hat{\mathbf{z}}_{0}^{i},\mathbf{H},\mathbf{X}^{i},\hat{\mathbf{X}}^{i})$triangleright model parameter gradient calculation
@@
22:              $\mathbf{z}_{T}^{i}\leftarrow\textsc{forward\_diffusion}(\mathbf{z}_{0}^{i},T)$triangleright maximum noising simulating inference time
23:              $\hat{\mathbf{z}}_{0}^{i}=\textsc{ddpm\_sample}(\mathcal{G}_{\Theta},\mathbf{z}_{T}^{i},T,\mathbf{H},c^{i})$triangleright full DDPM denoising loop
24:              $\hat{\mathbf{X}}=\textsc{sg}(\mathcal{D}(\mathbf{H},\hat{\mathbf{z}}_{0}^{i}))$triangleright decode predicted latent variable and stop gradient
@@
We use $w_{rec}=1$ and $w_{aux}=10000$ in our experiments. We do not use the SMPL losses $L_{SMPL}$ in training the denoiser model because the SMPL body model inference process slows down the training. The denoiser model is trained using an AdamW optimizer. The learning rate is set to $1e^{-4}$ with linear annealing. Our denoiser model is trained with scheduled training (Ling et al., 2020; Bengio et al., 2015), consisting of a fully supervised stage of 100K iterations, a scheduled stage of 100K iterations, and a rollout stage of 100K iterations. We set the maximum number of rollouts as 4. With the scheduled training, our latent motion primitive model can stably generate long motion sequences and better respond to the text prompt control even at poses that are not paired with the text prompt in the dataset.
@@ E.1 Experiment details
We apply DART to conduct online motion generation using the rollout algorithm 1 with a default seed motion $\mathbf{H}_{seed}$ of rest standing and using a classifier-free guidance weight of 5. We use the released checkpoints of FlowMDM (Barquero et al., 2024), TEACH (Athanasiou et al., 2022), and DoubleTake (Shafir et al., 2024) for baseline comparison. We adjust the handshake size and blending length of DoubleTake to be compatible with the shortest segment length of 15 frames. The history-conditioned modification of T2M-GPT is retrained on the BABEL dataset using the original hyperparameters. At generation time, the last frames of the previous action segment are encoded into tokens and provided as the first tokens when generating the next action segment to provide history conditioning of the previous action.
@@
E.2 Human Preference Studies.
We conduct human preference studies to quantitatively compare our method DART with baselines to provide a more comprehensive and convincing evaluation. We run human preference studies on Amazon Mechanical Turk(AMT) to evaluate generation realism and text-motion semantic alignment. Participants are given generation results from two different methods and are asked to select the generation that is perceptually more realistic or better aligns with the action text stream in subtitles, as illustrated in Fig. 4. During the realism evaluation, action descriptions were not displayed to eliminate distractions. We sample 256 sequences of action texts and durations from the BABEL dataset and use each method to generate motions given the action timelines. Participants are shown random pairs of results from our method and a baseline, and they are asked to choose the better generation. Each comparison is voted by 3 independent participants, with the video pairs randomly shuffled to ensure that participants do not know the method source.
Figure 4: Illustration of the human preference study interface for evaluating motion-text semantic alignment (top) and perceptual realism(bottom). Participants are requested to select the generation that is perceptually more realistic or better aligns with the action descriptions in subtitles (only visible in semantic preference study).
@@ Appendix F Latent diffusion noise optimization-based control
Optimization details. We use the Adam (Kingma & Ba, 2015) optimizer with a learning rate of 0.05 to optimize latent noises $\mathbf{Z}_{T}$ for 100 to 300 steps. The learning rate is linearly annealed to 0 and the gradient is normalized to stabilize optimization. The latent noises are initialized by sampling from Gaussian distributions. Empirically, we observe that initializing these latent noises with a small variance, such as 0.01, slightly improves the jerk and skate metrics at the cost of reduced motion diversity in the in-between experiment, compared to a unit variance initialization. The optimization running time is dependent on both the motion sequence length and the number of optimization steps. An example experiment of a target motion sequence of 60 frames and 100 optimization steps costs around 74 seconds.
Human-scene interaction. In human-scene interaction synthesis, we use two scene constraints to avoid human-scene collision and foot-floating artifacts. To reduce human-scene interpenetration, $cons\_coll(\cdot)$ penalizes joints that collide into scenes and exhibit negative SDF values in each frame:
$cons\_coll(\mathbf{J})=-\sum_{k=1}^{22}(\Psi(\mathbf{J}^{k})-\bm{\tau}^{k})_{-}$ (13)
where $\mathbf{J}^{k}$ denotes the location of the $k$-th joint in the scene coordinates, $\Psi(\cdot)$ denotes the signed distance function returning the signed distance from the query joint location to the scene, $\bm{\tau}$ denotes the joint-dependent contact threshold values, which are determined by the joint-skin distance in the rest pose, and $(\cdot)_{-}$ denotes clipping positive values. To reduce the occurrence of foot-floating artifacts, $cons\_cont(\cdot)$ encourages the foot to be in contact with the scene in every frame and is defined as:
$cons\_cont(\mathbf{J})=(\min_{k\in Foot}\Psi(\mathbf{J}^{k})-\tau)_{+}$ (14)
where $Foot$ denotes the set of foot joint indices, $\tau$ denotes the foot-floor contact threshold distance, and $(\cdot)_{+}$ denotes clipping negative values.
@@ Appendix G Reinforcement Learning-based Control
Architecture. Both the actor and critic networks are 4-layer MLPs with residual connections and a hidden dimension of 512. We apply the tanh scaling: $x=4\cdot tanh(x)$ (Ling et al., 2020) to the actor output to clip the action prediction in the range of $[-4,4]$, avoiding unbounded action predictions. The actor networks are initialized with close to zero weights to boost policy training following (Andrychowicz et al., 2021). The observation input contains a 512D text embedding, a 552D of history motion, a 1D observation of floor height relative to the human pelvis, a 1D observation of the floor plane distance from the pelvis to the goal location, and a 3D unit vector of the goal direction in the human-centric coordinates frames. We clamp the goal distance observation with a maximum value of 5m and the goal direction in the range of a 120-degree field of view to simulate egocentric human perception.
Rewards. The text-conditioned goal-reaching control policies are trained to maximize the discounted expectation of the following rewards (Zhang & Tang, 2022; Li et al., 2024a; Zhao et al., 2023):
@@
$r_{ori}=\frac{\langle\mathbf{p}^{i}-\mathbf{p}^{i-1},g-\mathbf{p}^{i-1}\rangle+1}{2},$ (17)
The moving orientation reward encourages the moving orientation to be aligned with the goal orientation, where $\mathbf{p}^{i}$ is the human pelvis location at step $i$, and $g$ is the goal location.
@@
The rewards are weighted with $w_{dist}=1$, $w_{succ}=1$, $w_{ori}=0.1$, $w_{skate}=100$, $w_{floor}=10$ for hopping and $w_{floor}=100$ for walking and running.
Appendix H Combination with physics-based motion tracking
Our method, DART, enables real-time motion generation in response to online text prompts. However, as a kinematic-based approach, DART may produce physically inaccurate motions with artifacts such as skating and floating. To address this, we demonstrate that DART can be integrated with physically simulated motion tracking methods, specifically PHC (Luo et al., 2023), to generate more physically plausible motions. In Fig. 5, we present an example sequence of a person crawling. The raw generation results from DART exhibit artifacts such as hand-floor penetration. Applying physics-based tracking to refine the raw motion successfully produces more physically plausible results, improving joint-floor contact and eliminating penetration artifacts. This integration combines the versatile text-driven motion generation of DART with the physical accuracy provided by the physics-based simulation. Given the real-time capabilities of both DART and PHC, it is possible to leverage physics to correct the kinematic motion generated by DART on the fly and then use the corrected motions for subsequent online generation.
(a) Crawling sequence generated by DART
(b) Physics-based motion tracking result
Figure 5: We demonstrate an example of integrating DART with the physics-based motion tracking method PHC (Luo et al., 2023) to achieve more physically plausible motions. The left image illustrates a crawling sequence generated by DART, exhibiting artifacts such as hand-floor penetration. The right image displays the physics-based motion tracking outcome applied to the raw generated sequence, which enhances joint-floor contact and resolves the hand-floor penetration issue.
Appendix I Discussion of long rollout results given a single text prompt
Our DART can autoregressively generate perpetual rollouts of actions that are inherently repeatable and extendable. For example, DART can produce minutes-long sequences of continuous human motion, such as jogging in circles, performing cartwheels, or dancing. These actions are inherently repeatable, and such extensions are also represented in the AMASS (Mahmood et al., 2019) dataset. DART can stably generate minutes-long rollouts of the same action, enabled by its autoregressive motion primitive modeling and scheduled training scheme.
Some other actions, however, have inherent boundary states that mark the completion of the action. For instance, “kneel down” reaches a boundary state where the knees achieve contact with the floor. Further extrapolation of “kneel down” beyond this boundary state is not represented in the dataset and is not intuitively anticipatable by humans, as no further motion logically extends within the action semantics. Continuing rollout using the “kneel down” text prompt results in motions exhibiting fluctuations around the boundary state.
In summary, long rollout results given a single text prompt will repeat naturally or fluctuate around a boundary state, depending on the inherent nature of the action and its representation in the dataset. We provide video results of minutes-long rollout generation results on our project website.
Appendix J Discussion on open-vocabulary motion generation
Limited vocabulary is a critical limitation and challenge shared by existing text-conditioned motion generation methods. Existing methods, including our approach DART, struggle to generalize to open-vocabulary text prompts due to the scarcity of 3D human motion data with text annotations. The scale of motion data available is several orders of magnitude smaller than that for text-conditioned image and video generation, primarily due to the reliance on marker-based motion capture systems, which are challenging to scale.
To expand the dataset and enable open-vocabulary generation, extracting human motion data from in-the-wild internet videos and generative image/video models (Kapon et al., 2024; Goel et al., 2023; Lin et al., 2023; Shan et al., 2024), is a promising direction. Additionally, the rapid advancement of vision-language models (VLMs) holds promise for automatically providing detailed, frame-aligned motion text labels to facilitate text-to-motion generation (Shan et al., 2024; Dai et al., 2023).
Appendix K Computing resources